OpenStudy (anonymous):

Could somebody explain how to do this: Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3> Answer choices: A. -9.1 Degrees B. 1.8 Degrees C. 0.9 Degrees D. 11.8 Degrees Not sure how to find an angle between vectors.

OpenStudy (anikhalder):

Do you know how to find the magnitude or length of a vector? And also so you know about scalar or dot product of 2 vectors?

OpenStudy (anonymous):

I know Scalar and Dot Product. @anikhalder I sort of know magnitude...

OpenStudy (anikhalder):

That's perfect! So let's go through this: What is the length of u vector?

OpenStudy (anonymous):

I'm not sure how to solve for length...

OpenStudy (anikhalder):

so its like, suppose i have a vector a = <3, 4> Then the length of the vector a is : \[\left| a \right|=\sqrt{3^{2} + {4^{2}}}\] \[\left| a \right|=\sqrt{25}\] That is the length of vector a is 5 So using the same way can you solve for the length of vector u?

OpenStudy (anonymous):

Oh! Is the magnitude |u| = SqRt. 41 ?

OpenStudy (anikhalder):

Perfect! You rock! What about the magnitude (or length) of vector v?

OpenStudy (anonymous):

|v| = 5 ?

OpenStudy (anikhalder):

Awesome! Now do you know the formula for dot (or scalar) product of 2 vectors a and b?

OpenStudy (anonymous):

Yes, would the product be: a x b = 32 ?

OpenStudy (anikhalder):

Yes that's correct! But, I will suggest when writing dot product don't use 'x' sign. You will learn that there is another type of product called the vector or cross product which uses this sign...but we don't need to worry about right now. Suppose I have a vector a and vector b Then the dot product of vectors a and b is also defined as: \[a.b=\left| a \right|\left| b \right|\cos \theta\] Where theta is the angle between the vectors a and b. So for our question at hand. Let's replace a by u and b by v. So, we get: \[u.v=\left| u\right|\left| v\right|\cos \theta\] We can replace u.v as 32 and the magnitudes of u and v on the right hand side of the equation to get: \[\cos \theta=\frac{ u.v }{ \left| u \right|\left| v \right|}\] Can you solve for costheta now?

OpenStudy (anonymous):

Is costheta = .99 ?

OpenStudy (anonymous):

@anikhalder

OpenStudy (anikhalder):

Yes! I find so as well. Now just find the cosine inverse of 0.99 and you'll get it in degrees

OpenStudy (anonymous):

I got 8.1 degrees, approximately, but that is not one of my answers.

OpenStudy (anikhalder):

That's what I get and now I must confess that I am confused as well

OpenStudy (anikhalder):

Let's verify

OpenStudy (anonymous):

Ok

OpenStudy (anikhalder):

http://onlinemschool.com/math/assistance/vector/angl/ Can you type the values in and see what you get for costheta because I get 0.99 in this website as well

OpenStudy (anonymous):

I am getting .99

OpenStudy (anonymous):

@anikhalder Can you help me with another problem, and then maybe go back to this one?

OpenStudy (anikhalder):

oh wait

OpenStudy (anonymous):

Express the complex number in trigonometric form. -3 + 3 square root of three i

OpenStudy (anonymous):

And I'm confused as to what that calculator meant.

OpenStudy (anikhalder):

Check this, we were just using 0.99, we cant do that we have to use the full result which was 0.9995120761 plug this and find the cos inverse. You will see that we get our answer :)

OpenStudy (anonymous):

Oh. I see. So the answer would be 1.8 degrees?

OpenStudy (anikhalder):

Yep! And yes just take 4 minutes and see this video on complex numbers! he explains better than me :)))

OpenStudy (anonymous):

What video?

OpenStudy (anikhalder):

Sorry...my bad i forgot to give the link: https://www.youtube.com/watch?v=6z6fzPXUbSQ

OpenStudy (anonymous):

Thank you so much. (: I watched that video, though, and am still confused on my example. He helped me with a couple others, but the one I posted about is a bit confusing after watching the video.

OpenStudy (anikhalder):

You mean its like : \[3+3\sqrt{3}i\] So, \[\left| z \right| = \sqrt{3^{2}+(3\sqrt{3})^{2}} = 6\] and \[\tan \theta = \frac{ 3\sqrt{3} }{ 3}\] i.e. \[\theta = 60 degrees\] Now, in polar or trigonometric form, z = \[\left| z \right|(\cos \theta + i \sin \theta)\] Just substitute the values and find the answer :)))

OpenStudy (anonymous):

What is the Z value?

OpenStudy (anonymous):

And is the theta value 120 degrees?

OpenStudy (anikhalder):

you can say \[\left| z \right|\] is like the absolute value of the complex number (like the